OpenStudy (anonymous):

If y=1/4 when x=1/8, find x when y=3/16. Determine the Solution.

OpenStudy (anonymous):

y=1/4 y=2x x=1/8 x=1/2y y=3/16 x=1/2(3/16) x=3/32